OpenStudy (anonymous):

What is csc(9pi/4) ?

OpenStudy (anonymous):

well uh..Pi radians means 180 degrees, so in degrees this would simplify to - csc(9x180/4) = csc(9x45) = csc 405 degrees. Now cosec 360 is nothing but cosec 0 (since 0 and 360 are actually the same angle, just that 360 is one complete revolution.) so you're left with the next 45 degrees. (360+45 = 405) so the value would be cosec 45.